OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -one fourthx - 5. y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (johnweldon1993):

perpendicular means the slope is the negative reciprocal of the other line... so this lines slope is -1/4x the negative reciprocal is 4/1 or just 4x now use point slope again y - 7 = 4(x -(-1) y - 7 = 4(x+1) y - 7 = 4x + 4 y = 4x + 11 :)